# Title keyword 5 for citation Title for citation Title for citation

### v2

@@ 5 Wireless Network Large Model
As technology advances, large-scale AI models (LAM) exhibit remarkable capabilities and extensive potential in theoretical research and practical applications. However, this rapid development trend presents new challenges and requirements for the supporting network infrastructure. As a crucial component of modern communication infrastructure, wireless networks directly impact large-scale AI models’ data transmission efficiency, processing speed, and overall performance. Therefore, this chapter aims to explore the critical role of wireless networks in supporting the operation of these models and conduct a systematic analysis and reflection on the relevant issues related to wireless networks in the context of constructing large-scale AI models.
5.1 Differences between LLMs and Wireless Network Large Model
As a token-based approach, LLM[261] has demonstrated outstanding performance in text and image processing tasks. Compared to the standardized text and image inputs typically handled by LLMs, the data from wireless communication systems covering air interface protocols, core network architectures, and operational management present significant complexity and diversity. Wireless communication data encompasses standardized and non-standardized heterogeneous data types characterized by temporal and relational features.
@@ 5.1 Differences between LLMs and Wireless Network Large Model
Specifically, the air interface involves wireless channels’ transmission and scheduling management, encompassing diverse information generated by various wireless communication standards, such as LTE and 5G. The core network handles routing, authentication, and service control of user data, which includes sensitive content such as user behavior and location information, demanding high levels of privacy and security. Operational management data covers status monitoring, fault diagnosis, and performance optimization of network equipment. These datasets are typically large-scale, high-dimensional, lacking effective designs for unifying data across different standards. These characteristics necessitate the consideration of complex factors such as spatiotemporal relationships, real-time requirements, and system stability when processing and analyzing wireless communication data, presenting significant differences and challenges compared to traditional LLM data processing approaches, as illustrated in Figure 23.
Figure 23: Differences between LLMs and wireless network large model
@@
Moreover, wireless communication systems constitute a highly complex ecosystem[262] employing multiple technological paradigms to meet diverse technical requirements and environmental scenarios. For instance, coding and modulation techniques are pivotal in ensuring data transmission reliability and efficiency; MIMO technology enhances data transfer rates and signal quality by utilizing multiple transmit and receive antennas; wireless interference management aims to increase overall channel capacity; traffic forecasting facilitates effective network planning and resource allocation; load balancing techniques optimize the distribution of user and data traffic across different network resources; and IP route optimization maximizes the efficiency of data packet transmission paths. These disparate wireless task models exhibit significant differences, and their commonalities are not always clear, contributing to the need for a nuanced approach to developing a large wireless network model that diverges from traditional LLM methodologies.
@@
In contrast to the relatively lenient requirements for inference speed in LLM models used in text and image processing, the demands placed on model inference speed in wireless communication tasks are critically important. Insufficient inference speed in a large wireless network model can lead to outdated results due to significant changes in channel conditions and signal characteristics during the inference process. Therefore, ensuring that models in wireless communication systems can quickly obtain inference results is a key performance indicator for deployment effectiveness.
@@
On the other hand, the requirements for inference accuracy in large wireless network Models far exceed those of LLMs. In language models, different combinations of word arrangements have minimal impact on semantic expression; however, in wireless communication systems, which rely on exact mathematical models, sensitivity to inference errors is markedly high, with repercussions that are difficult to quantify. Signal modulation, a critical process encoding information onto carrier signals, exemplifies this sensitivity. Minor deviations in predicting modulation parameters by the model, such as errors in frequency or phase, can prevent the receiver from correctly demodulating signals, resulting in data loss or errors. In distributed wireless networks, even slight errors in predicting clock offsets or sync signals by the model can lead to synchronization failures among nodes, thereby compromising network stability. Therefore, precise control over the inference accuracy of the Wireless Network Large Model is a critical research challenge.
@@
5.2 Potential Approaches to Constructing Wireless Network Large Model
Wireless Network Large Model[263] refers to an intelligent and efficient solution for communication network operations, achieved through integrating knowledge and technologies from the communication domain. The Wireless Network Large Model provides network management and operations fault detection and diagnostic services. Traditional network faults often lead to performance degradation and service interruptions. By analyzing massive network data, the Wireless Network Large Model can monitor network status in real time, promptly identify and locate potential fault points, and provide accurate fault diagnosis and handling recommendations to operations personnel.
@@ 5.2 Potential Approaches to Constructing Wireless Network Large Model
Figure 24: The three-layer architecture of Wireless Network Large Model
@@
The prevailing industry viewpoint suggests that the wireless network large model should adopt a three-layer model structure, as demonstrated in Figure 24. The L2 layer can provide customized solutions tailored to specific business scenarios in this framework. For instance, the L2 layer model can optimize energy utilization efficiency in energy conservation within networks by analyzing data traffic and device loads. Regarding traffic prediction and fault monitoring, the L2 layer model can employ deep learning techniques to identify anomalies and issue early warnings, thereby enhancing network stability and reliability.
Subsequently, the L1 layer involves domain-specific grand models. The design of the L1 layer emphasizes modeling and optimization for specific domain challenges and issues. For example, the L1 layer model can analyze wireless channel conditions and user behavior in the air interface to improve data transmission rates and coverage. The L1 layer model can utilize network topology and device status information for fault diagnosis and optimization scheduling in the core network and operations. Lastly, the L0 layer represents the universal grand model for the entire network. The design of this layer needs to consider the overall architecture and operational mechanisms of communication systems to ensure it possesses strong generalization capabilities across various scenarios. Through comprehensive data analysis and intelligent decision-making, it aims to provide optimization and management services for all communication tasks.
@@
Firstly, considering the architecture design of the Wireless Network Large Model in communication networks, constructing a universally applicable large model for the entire network faces numerous challenges. On the one hand, data within communication networks exhibits highly structured characteristics, encompassing extensive information from various devices and sensors[266], including network traffic, device status, and user behavior. The diversity and complexity of these data present significant challenges for model construction and optimization. On the other hand, tasks within communication networks exhibit complex interdependencies and relationships, with difficulties in clearly defining their commonalities and differences. For instance, network traffic characteristics may vary substantially across different business scenarios, rendering traditional unified modeling approaches less suitable. Therefore, directly constructing a universally applicable large-scale model for the entire network is exceptionally challenging.
A phased approach to model construction becomes particularly crucial to address the aforementioned challenges effectively. This strategy reduces the complexity of model design and optimization and enhances scalability and adaptability, providing reliable foundational support for developing and optimizing large-scale network models. When implementing the phased exploration strategy, the initial focus should be small-scale, offline L2-level models. Through the in-depth research and practical validation of critical business scenarios such as network energy efficiency, traffic prediction, and fault monitoring, efficient utilization of network resources can be achieved, thereby improving overall network performance and sustainability. Subsequently, progression towards large-scale, real-time responsive L1 and L0-level models should be undertaken. In designing L1-level models, emphasis should be placed on addressing domain-specific challenges and issues, employing targeted modeling and optimization approaches to achieve domain generalization capabilities across air interface, operations, and core network sides. To construct a universally applicable L0-level model encompassing the entire network, carefully considering the communication system’s overall architecture and operational mechanisms is essential to ensure the model’s ability to support services and strategies for all communication tasks. Constructing datasets is also crucial for large-scale network models[267] directly impacting AI models’ training efficacy and application performance. Current challenges primarily revolve around data acquisition and quality difficulties [268]. Addressing these issues requires effective measures to ensure dataset quality and completeness.
@@
Regarding the computational demands of 6G communication to support large-scale network models, optimization efforts are needed in both AI training and deployment. This aims to achieve efficient, reliable, secure, and stable data transmission and utilization of computing resources. In AI training services[271], key factors include efficient and reliable training data collection and transmission processes, proprietary protocol design, support from network elements, and customized traffic scheduling algorithms. Proprietary protocols tailored explicitly for AI task transmission should be developed in the design of access and core networks in the design of access and core networks. These protocols need to account for the unique characteristics of AI tasks, such as large data volumes and computational intensity, to ensure optimal transmission efficiency and performance. Secondly, critical network elements must fully support AI task transmission, encompassing network and computing devices. These devices should have sufficient computing and storage capabilities to handle the extensive data and complex computational demands generated during AI training.
@@
Traditional task scheduling algorithms may fail to adequately account for load conditions and computational resource utilization among nodes, resulting in decreased system performance or uneven workload distribution. Hence, intelligent task scheduling algorithms must be designed to dynamically monitor and analyze load conditions across nodes and evaluate inference task priorities and computational resource requirements in real time. This approach allows for the rational allocation and scheduling of inference tasks, effectively enhancing overall system performance and ensuring optimal utilization of node resources to achieve efficient completion of inference tasks.
@@
5.3 Challenges of Wireless Network Large Model
@@
Data Heterogeneity: In wireless networks, the diversity and complexity of data present significant challenges for constructing adequate large-scale model datasets and inference inputs. These data encompass a wide range of types[274], including uplink and downlink channel information, Measurement Reports, Minimization of Drive Test data, Key Performance Indicator (KPI) monitoring data, network topology performance metrics, as well as cost and energy consumption data, among various standardized or non-standardized formats. These data may exhibit significant granularity and temporal resolution differences from diverse sources. A primary research concern is effectively aligning such heterogeneous data to establish a consistent and representative dataset. The differences in data formats, sources, and collection environments must be carefully considered during data preprocessing to ensure high accuracy and reliability in training and inference for the final model. Techniques such as data normalization, time-series alignment, data fusion, and handling of missing data may be necessary. Furthermore, for constructing a large wireless network model, attention must be given to converting processed data into suitable input formats for inference algorithms. This may involve feature extraction and selection, data dimensionality reduction, and constructing appropriate input representations to enable efficient inference and decision-making by the model.
@@
Real-Time Inference: Wireless Network Large Model is commonly employed for real-time decision-making and optimization tasks[275] such as dynamic spectrum allocation, power control, and user access management. However, if the inference time of these models is prolonged or inefficient, timely responses to network changes and demands may be compromised[276], thereby impacting overall network performance and user experience. Consequently, deploying the large wireless network model demands exceptionally high performance standards. This is particularly critical in network resource scheduling, where stringent time constraints necessitate model inference completion within milliseconds to effectively manage and allocate communication network resources. This directly influences the network’s real-time responsiveness and service quality. To meet these requirements[277], the design and deployment of the Wireless Network Large Model must prioritize optimizing inference speed. This involves hardware acceleration, algorithm optimization, and adjustments to network architecture[278]. Moreover, achieving a balance between model accuracy and computational efficiency is essential to enable efficient and reliable management and scheduling of network resources.
@@
Ambiguous Route: In wireless communication technology, constructing a broadly applicable large-model framework continues to present challenges. At the same time, current industry practices favor leveraging large-scale pre-trained language models as foundational elements, subsequently fine-tuned for specific industry applications. The complexity and dynamics of wireless networks demand models with heightened adaptability and flexibility. Recent explorations in semantic communication-based models[279] have made strides towards a more profound understanding and processing of data flows within wireless networks. However, the feasibility and efficacy of these models in practical deployment, including their stability and performance across diverse environments and conditions, remain open research questions requiring further empirical validation.
@@
Security Threats: Wireless Network Large Model faces multiple security challenges[280]. Firstly, there are concerns regarding data privacy and protection. In deploying 5G networks, large datasets may include sensitive information such as user locations, mobility patterns, and communication content. Without adequate protection, these data could risk privacy breaches[281], leading to compromised user trust and legal issues. Secondly, data tampering presents another significant threat to large models. Attackers may inject misleading data or manipulate sources to influence the model’s learning process and decision outcomes. For instance, incorrect input data could misjudge user demands or network congestion conditions. Additionally, network attacks targeting large wireless network models are hazardous. Malicious actors may disrupt model operations or regular requests through flooding or resource consumption attacks, severely impacting service availability and real-time decision-making capabilities. Finally, ensuring the security and integrity of large model architectures is crucial. Reverse engineering and leakage of model parameters could provide attackers with sensitive information about intellectual property or operational mechanisms, jeopardizing service stability and reliability.
@@ 6 AI as a Service
As described in Section 1, to more effectively support ”native intelligence” and achieve ubiquitous ”universal intelligence,” future 6G networks will treat AI as a service, giving rise to the concept of 6G AI as a Service (AIaaS) [282], which is presented in Figure 26. 6G will provide efficient end-to-end support for AI-related businesses and applications, intelligently connecting distributed agents for large-scale AI deployment across various industries. 6G AI services will address needs for high real-time performance, high security and privacy, or low overall energy consumption by conducting AI training or inference within the network, offering intelligent capabilities tailored to different application scenarios.
6G networks, as inherently intelligent architectures, will be able to efficiently train or infer large-scale distributed AI through resources and functionalities within the network, including communication, computation, datasets, and foundational models [283]. This transition from AI for networks in 5G to networks for AI in 6G will ultimately provide intelligence that goes beyond data, directly targeting users. This transformation will redefine the ecosystem of edge clouds and create new business models through 6G mobile networks, transitioning from connectivity-focused to service-oriented networks and ultimately achieving ubiquitous intelligence.

### v3

@@ 4 Wireless Network Large Model
Large-scale AI models (LAM) exhibit remarkable capabilities and great potential in academia and industries. However, such a trend presents new challenges for the underlying network infrastructure. Wireless networks play a vital role in the infrastructure and are directly associated with LAM’s performance, such as data transmission, processing speed, etc. Therefore, this chapter aims to explore the critical role of wireless networks in supporting the operation of these LAM and conduct a systematic analysis and reflection on the relevant issues related to wireless networks in the context of constructing LAM.
4.1 Comparisons Between LLM and Wireless Network Large Model
LLMs[227] have demonstrated their promise in text and image tasks. It is challenging to directly apply traditional LLMs to wireless communication systems, as the data in the system consists of various protocol data in access networks, CNs, and application data from the upper layer. Furthermore, these data are both structured and non-structured and may be heterogeneous with temporal and relational features.
@@ 4.1 Comparisons Between LLM and Wireless Network Large Model
Specifically, the air interface involves wireless channels’ transmission and scheduling management, encompassing diverse information generated by various wireless communication standards, such as LTE and 5G. The CN handles routing, authentication, and service control of user data, which includes sensitive content such as user behavior and location information, demanding high levels of privacy and security. Operational management data covers status monitoring, fault diagnosis, and performance optimization of network equipment. These datasets are typically large-scale, high-dimensional, lacking effective designs for unifying data across different standards. These characteristics necessitate the consideration of complex factors such as spatiotemporal relationships, real-time requirements, and system stability when processing and analyzing wireless communication data, presenting significant differences and challenges compared to conventional LLM data processing approaches, as illustrated in Figure 22.
Figure 22: Differences between LLM and wireless network large model
@@
Moreover, wireless communication systems constitute a highly complex ecosystem[228] that employs multiple technological paradigms to meet diverse technical requirements and environmental scenarios. For example, coding and modulation techniques are pivotal in ensuring data transmission reliability and efficiency; MIMO technology enhances data transfer rates and signal quality by utilizing multiple transmit and receive antennas; wireless interference management aims to increase overall channel capacity; traffic forecasting facilitates effective network planning and resource allocation; load balancing techniques optimize the distribution of user and data traffic across different network resources; and internet protocol (IP) route optimization maximizes the efficiency of data packet transmission paths. These disparate wireless task models exhibit significant differences, and their commonalities are not always clear, contributing to the need for a nuanced approach to developing a large wireless network model that diverges from conventional LLM methodologies.
@@
In contrast to the relatively lenient requirements for inference speed in LLM models used in text and image processing, the demands placed on model inference speed in wireless communication tasks are critically important. Insufficient inference speed in a large wireless network model can lead to outdated results due to significant changes in channel conditions and signal characteristics during the inference process. Therefore, ensuring that models in wireless communication systems can quickly obtain inference results is a KPI for deployment effectiveness.
@@
On the other hand, the requirements for inference accuracy in large wireless network models far exceed those of LLM. In language models, different combinations of word arrangements have minimal impact on semantic expression; however, in wireless communication systems, which rely on exact mathematical models, sensitivity to inference errors is markedly high, with repercussions that are difficult to quantify. Signal modulation, a critical process encoding information onto carrier signals, exemplifies this sensitivity. Minor deviations in predicting modulation parameters by the model, such as errors in frequency or phase, can prevent the receiver from correctly demodulating signals, resulting in data loss or errors. In distributed wireless networks, even slight errors in predicting clock offsets or sync signals by the model can lead to synchronization failures among nodes, thereby compromising network stability. Therefore, precise control over the inference accuracy of the wireless network large model is a critical research challenge.
@@
4.2 Potential Approaches to Constructing a Wireless Network Large Model
Wireless network large model[229] refers to an intelligent and efficient solution for communication network operations, achieved through integrating knowledge and technologies from the communication domain. The wireless network large model provides network management and operations fault detection and diagnostic services. The failures in networks may lead to performance degradation and service interruptions. By analyzing massive network data, the wireless network large model can monitor network status in a real-time fashion, quickly identify and locate potential reasons, and provide accurate diagnosis strategies.
@@ 4.2 Potential Approaches to Constructing a Wireless Network Large Model
The prevailing industry viewpoint suggests that the wireless network large model should adopt a three-layer model structure, as demonstrated in Figure 23. The L2 layer can provide customized solutions tailored to specific business scenarios in this framework. For instance, the L2 layer model can optimize energy utilization efficiency in energy conservation within networks by analyzing data traffic and device loads. Regarding traffic prediction and fault monitoring, the L2 layer model can employ DL techniques to identify anomalies and issue early warnings, thereby enhancing network stability and reliability.
Subsequently, the L1 layer involves domain-specific grand models. The design of the L1 layer emphasizes modeling and optimization for specific domain challenges and issues. For example, the L1 layer model can analyze wireless channel conditions and user behavior in the air interface to improve data transmission rates and coverage. The L1 layer model can utilize network topology and device status information for fault diagnosis and optimization scheduling in the CN and operations. Lastly, the L0 layer represents the universal grand model for the entire network. The design of this layer needs to consider the overall architecture and operational mechanisms of communication systems to ensure it possesses strong generalization capabilities across various scenarios. Through comprehensive data analysis and intelligent decision-making, it aims to provide optimization and management services for all communication tasks.
Figure 23: The three-layer architecture of a wireless network large model
@@
Firstly, considering the architecture design of the Wireless Network Large Model in communication networks, constructing a universally applicable large model for the entire network faces numerous challenges. On the one hand, data within communication networks exhibits highly structured characteristics, encompassing extensive information from various devices and sensors[232], including network traffic, device status, and user behavior. The diversity and complexity of these data present significant challenges for model construction and optimization. On the other hand, tasks within communication networks exhibit complex interdependencies and relationships, with difficulties in clearly defining their commonalities and differences. For instance, network traffic characteristics may vary substantially across different business scenarios, rendering conventional unified modeling approaches less suitable. Therefore, directly constructing a universally applicable large-scale model for the entire network is exceptionally challenging.
A phased approach to model construction becomes particularly crucial to address the aforementioned challenges effectively. This strategy reduces the complexity of model design and optimization and enhances scalability and adaptability, providing reliable foundational support for developing and optimizing large-scale network models. When implementing the phased exploration strategy, the initial focus should be small-scale, offline L2-level models. Through the in-depth research and practical validation of critical business scenarios such as network energy efficiency, traffic prediction, and fault monitoring, efficient utilization of network resources can be achieved, thereby improving overall network performance and sustainability. Subsequently, progression towards large-scale, real-time responsive L1 and L0-level models should be undertaken. In designing L1-level models, emphasis should be placed on addressing domain-specific challenges and issues, employing targeted modeling and optimization approaches to achieve domain generalization capabilities across air interface, operations, and CN sides. To construct a universally applicable L0-level model encompassing the entire network, carefully considering the communication system’s overall architecture and operational mechanisms is essential to ensure the model’s ability to support services and strategies for all communication tasks. Constructing datasets is also crucial for large-scale network models[233] directly impacting AI models’ training efficacy and application performance. Current challenges primarily revolve around data acquisition and quality difficulties [234]. Addressing these issues requires effective measures to ensure dataset quality and completeness.
@@
Regarding the computational demands of 6G communication to support large-scale network models, optimization efforts are needed in both AI training and deployment. This aims to achieve efficient, reliable, secure, and stable data transmission and utilization of computing resources. In AI training services[237], key factors include efficient and reliable training data collection and transmission processes, proprietary protocol design, support from network elements, and customized traffic scheduling algorithms. Proprietary protocols tailored explicitly for AI task transmission should be developed in the design of access and CNs in the design of access and CNs. These protocols need to account for the unique characteristics of AI tasks, such as large data volumes and computational intensity, to ensure optimal transmission efficiency and performance. Secondly, critical network elements must fully support AI task transmission, encompassing network and computing devices. These devices should have sufficient computing and storage capabilities to handle the extensive data and complex computational demands generated during AI training.
@@
conventional task scheduling algorithms may fail to adequately account for load conditions and computational resource utilization among nodes, resulting in decreased system performance or uneven workload distribution. Hence, intelligent task scheduling algorithms must be designed to dynamically monitor and analyze load conditions across nodes and evaluate inference task priorities and computational resource requirements in real time. This approach allows for the rational allocation and scheduling of inference tasks, effectively enhancing overall system performance and ensuring optimal utilization of node resources to achieve efficient completion of inference tasks.
@@
4.3 Challenges of Deploying Wireless Network Large Model
@@
Data Heterogeneity: In wireless networks, the diversity and complexity of data present significant challenges in constructing adequate large-scale training and testing corpus. There exists a wide range of structured and unstructured [240] data in networks. These data include uplink and downlink channel information, measurement reports, Minimization of Drive Test data, KPI monitoring data, network topology performance metrics, as well as cost and energy consumption logs. They are highly heterogeneous and exhibit significant differences in granularity and temporal resolution from diverse sources. A primary research concern is effectively aligning such heterogeneous data to establish a consistent and representative dataset. Exploring techniques such as data normalization, time-series alignment, and data fusion can help to build high-quality training corpus. Furthermore, for constructing a large wireless network model, attention must be given to converting processed data into suitable input formats for inference algorithms. This may involve feature extraction and selection, data dimensionality reduction, and constructing appropriate input representations to enable efficient inference and decision-making by the model.
@@
Real-Time Inference: Wireless network large model can be employed for real-time decision-making tasks[241] such as dynamic spectrum allocation, power control, and user access management. However, the long inference time of these large models [242] may degrade overall network performance and user experience. This is particularly critical in network resource scheduling, where stringent time constraints necessitate model inference completion within milliseconds to effectively manage and allocate communication network resources. This directly influences the network’s real-time responsiveness and service quality. To meet these requirements[243], the design and deployment of the wireless network large model must prioritize optimizing inference speed. This involves hardware acceleration, algorithm optimization, and adjustments to network architecture[244]. Moreover, achieving a balance between model accuracy and computational efficiency is essential to enable efficient and reliable management and scheduling of network resources.
@@
Ambiguous Route: It is challenging to build a generalized large-model framework for wireless communications. At the same time, current industry practices favor leveraging large-scale pre-trained language models as foundational elements, subsequently fine-tuned for specific industry applications. The complexity and dynamics of wireless networks demand models with heightened adaptability and flexibility. Recent explorations in semantic communication-based models[245] have made strides towards a more profound understanding and processing of data flows within wireless networks. However, the feasibility and efficacy of these models in practical deployment, including their stability and performance across diverse environments and conditions, remain open research questions requiring further empirical validation.
@@
Security Threats: Wireless network large model faces multiple security challenges [246]. Firstly, these models may suffer from privacy attacks. Datasets collected from 5G networks may involve sensitive information such as user locations, mobility patterns, and communication content. Without adequate protection, these data face privacy threats [247], leading to compromised user trust and legal issues. Secondly, data tampering presents another significant threat to large models. Attackers may inject misleading data or manipulate sources to influence the model’s learning process and decision outcomes. For instance, incorrect input data could misjudge user demands or network congestion conditions. Additionally, network attacks targeting large wireless network models are hazardous. Malicious actors may disrupt model operations or regular requests through flooding or resource consumption attacks, severely impacting service availability and real-time decision-making capabilities. Finally, ensuring the security and integrity of large model architectures is crucial. Reverse engineering and leakage of model parameters could provide attackers with sensitive information about intellectual property or operational mechanisms, jeopardizing service stability and reliability.
@@ 5 AI as a Service
As described in Section 1, to more effectively support “native intelligence” and achieve ubiquitous “universal intelligence”, future 6G networks will treat AIaaS, giving rise to the concept of 6G AIaaS [248], which is presented in Figure 25. 6G will provide efficient end-to-end support for AI-related businesses and applications, intelligently connecting distributed agents for large-scale AI deployment across various industries. 6G AI services will address needs for high real-time performance, high security and privacy, or low overall energy consumption by conducting AI training or inference within the network, offering intelligent capabilities tailored to different application scenarios.
6G networks, as inherently intelligent architectures, will be able to efficiently train or infer large-scale distributed AI through resources and functionalities within the network, including communication, computation, datasets, and foundational models [249]. This transition from AI4NET in 5G to networks for AI in 6G will ultimately provide intelligence that goes beyond data, directly targeting users. This transformation will redefine the ecosystem of edge clouds and create new business models through 6G mobile networks, transitioning from connectivity-focused to service-oriented networks and ultimately achieving ubiquitous intelligence.